OpenStudy (anonymous):

Solve: 6(2x - 1) = -42 A. x = -4 B. x = 3 C. x = -3.42 D. x = -3 Solve: 2(x + 5) - 4(3x + 2) = -48 A. x = -5 B. x = 5 C. x = 5.5 D. x = 4.2

OpenStudy (anonymous):

im trying to figure it out.

OpenStudy (anonymous):

1.) D 2.) B

OpenStudy (merchandize):

6(2x-1)=-42 12x-6=-42 12x=-42+6 12x=-36 x=-36/12 x=-3

OpenStudy (anonymous):

i got the first one figuredout im in the middle of the second one now.

OpenStudy (merchandize):

2(x+5)-4(3x+2)=-48 2x+10-12x-8=-48 -10x+2=-48 -10x=-48-2 -10x=-50 x=50/10 x=5

OpenStudy (anonymous):

thank i have my answers on paper i have to scan it and turn my work .

OpenStudy (anonymous):

thank you i got stuck on 2 and you helped me i figured out where i went wrong

OpenStudy (anonymous):

thank everyone.